# Resilience for Regular Path Queries: Towards a Complexity Classification

### v2

@@ 6. Hardness of Finite Languages Having Repeated Letters
First, we must have $\gamma_{1}\neq\epsilon$, otherwise $\alpha^{\prime}$ is an infix of $a\gamma$ which is a strict infix of $a\gamma a\delta\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is reduced.
@@
Hence, applying Section 6.1 to $L^{\prime}\coloneq L$, we know that $L$ contains a maximal-gap word $\alpha=a\gamma a\delta$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ with $\gamma_{1}$ and $\gamma_{2}$ being respectively a non-empty suffix of $\gamma$ and a non-empty prefix of $\gamma$. We now claim that, if $\delta\neq\epsilon$, then $L$ is four-legged, allowing us to conclude:
{claimrep}
Let $L^{\prime}$ be a reduced language containing a word $a\gamma a\delta$ such that $\delta\neq\epsilon$ and a word $\gamma_{1}a\gamma_{2}$ such that $\gamma_{1}\neq\epsilon$, $\gamma_{2}\neq\epsilon$, and $\gamma_{1}$ is a suffix of $\gamma$. Then $L^{\prime}$ is four-legged.
@@
Indeed, the word $\alpha^{\prime\prime}$ is an infix of $\gamma a\delta$ (because $\gamma_{1}$ is a suffix of $\gamma$), which is a strict infix of $a\gamma a\delta$ which is in $L^{\prime}$. Thus, as $L^{\prime}$ is reduced, $\alpha^{\prime\prime}\notin L^{\prime}$, so we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
Applying Section 6.1 with $L^{\prime}\coloneq L$, we can conclude unless $\delta=\epsilon$, which we assume in the rest of the proof.
@@
The overlapping case is when we have $|\gamma_{1}|+|\gamma_{2}|>|\gamma|$, so that we can write $\gamma=\eta^{\prime\prime}\eta\eta^{\prime}$ such that $\gamma_{1}=\eta\eta^{\prime}$ and $\gamma_{2}=\eta^{\prime\prime}\eta$, with the “overlap” $\eta$ being nonempty.
@@
Let us first deal with the overlapping case. We have $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, meaning $\alpha^{\prime}=\eta\eta^{\prime}a\eta^{\prime\prime}\eta\in L$, and we know that the “overlap” $\eta$ is nonempty. We then show:
{claim}
@@
Let us first assume that $\eta^{\prime}$ is nonempty and show that $L$ is four-legged in this case. Let us write $\eta^{\prime}=x\sigma$. Substituting this into $\alpha^{\prime}$ and $\alpha$, we have $\alpha^{\prime}=\eta x(\sigma a\eta^{\prime\prime}\eta)\in L$ and $\alpha=(a\eta^{\prime\prime}\eta)x(\sigma a)\in L$. Now, taking $x$ as body, let us consider the cross-product word $\alpha^{\prime\prime}=\eta x\sigma a=\eta\eta^{\prime}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta\eta^{\prime}a\in L$. Hence, as $L$ is reduced, $\alpha^{\prime\prime}\notin L$, so that we can conclude that $L$ is four-legged unless $\eta^{\prime}$ is empty.
Let us substitute $\eta^{\prime}=\epsilon$ in $\alpha$ and $\alpha^{\prime}$. We know that $\alpha^{\prime}=\eta a\eta^{\prime\prime}\eta\in L$ and $\alpha=a\eta^{\prime\prime}\eta a\in L$. Let us now assume that $\eta^{\prime\prime}$ is nonempty and show that $L$ is four-legged in that case as well. Let us write $\eta^{\prime\prime}=x\sigma$. We have $\alpha=ax(\sigma\eta a)\in L$ and $\alpha^{\prime}=(\eta a)x(\sigma\eta)\in L$. Now, taking $x$ as body again, let us consider the cross-product word $\alpha^{\prime\prime}=ax\sigma\eta=a\eta^{\prime\prime}\eta$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta a\in L$. Hence, as $L$ is reduced, $\alpha^{\prime\prime}\notin L$, so we can conclude again that $L$ is four-legged unless $\eta^{\prime\prime}$ is empty.
@@
From Section 6.1, we can conclude unless $\eta^{\prime}=\eta^{\prime\prime}=\epsilon$, so we assume this in the rest of the proof of hardness for the overlapping case. We now know that $\alpha^{\prime}=\eta a\eta\in L$ and $\alpha=a\eta a\in L$.
Let us now re-use the fact that $\alpha$ was chosen to be a maximal-gap word of $L$. Now, as $\eta$ is non-empty, observe that $\alpha^{\prime}=\eta a\eta$ contains repeated letters across the two copies of $\eta$. These letters are spaced at least by the length of $\eta$ (and possibly more, if $\eta$ contains repeated letters – even though this cannot happen by the maximal-gap condition). In any case, $\alpha^{\prime}$ achieves a gap between two repeated letters which is at least that of $\alpha$: so $\alpha$ and $\alpha^{\prime}$ are tied for the first criterion in the definition of maximal-gap. Hence, by the second criterion, since $\alpha$ is maximal-gap, we know that $\alpha^{\prime}$ cannot be strictly longer than $\alpha$. So from $|\alpha^{\prime}|\leq|\alpha|$, as $|\alpha^{\prime}|=2|\eta|+1$ and $|\alpha|=|\eta|+2$, we conclude $|\eta|\leq 1$. As $\eta$ is nonempty, we have $|\eta|=1$. This letter may be different from $a$ or not. So we can conclude the overlapping case by showing two hardness results:
{claim}
Let $L^{\prime}$ be any reduced language containing $aba$ and $bab$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is hard.
Figure 10. Gadget used in the proof of Section 6.1
@@
We use the gadget whose completion is pictured on Figure 10 (left). It is obviously a pre-gadget, and one can check that the hypergraph of matches can be reduced to the odd chain of Figure 10 (right), so that it is indeed a gadget and we can conclude by Section 4.8. More precisely, observe that there are three kinds of directed paths in the gadget:
  • Paths of length 2 or less: all of these form a word which is a strict infix of $aba$ or $bab$, so they are not in the language by reducedness.
  • Paths of length 3: all of these form the word $aba$ or $bab$, so they are in the language.
@@
[…]
{claim}
Let $L$ be any reduced language containing the word $aaa$. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
Figure 11. Gadget used in the proof of Section 6.1
@@
We use the gadget whose completion is pictured on Figure 11 (left) (which incidentally turns out to be identical to that of Figure 3(a) in Proposition 4.1). This is indeed a pre-gadget, and the hypergraph of matches can be condensed to Figure 11 (right), so it is indeed a gadget. Note that we use the fact that all directed paths in the gadget form words of the form $a^{*}$: by reducedness of the language, the matches precisely correspond to the words of length 3. ∎
@@
We conclude the proof of Section 6.1 by covering the non-overlapping case. Recall that, in this case, we know that the language $L$ contains a maximal-gap word $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, and recall that we already know that $\gamma_{1}$ and $\gamma_{2}$ are both non-empty.
@@
As $\gamma_{1}$ is nonempty, let us write $\gamma_{1}=\chi^{\prime}x$. If $\gamma_{1}$ has length 2 or more, then $\chi^{\prime}$ is nonempty, and we have $\alpha^{\prime}=\chi^{\prime}x(a\gamma_{2})\in L$ and $\alpha=(a\gamma_{2}\eta\chi^{\prime})xa\in L$. Let us take $x$ as body and consider the cross-product word $\alpha^{\prime\prime}=\chi^{\prime}xa=\gamma_{1}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$. So, as $L$ is reduced, we have $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged. Hence, it suffices to focus on the case where $\gamma_{1}=x$ has length 1.
Now, as $\gamma_{2}$ is nonempty, let us write $\gamma_{2}=y\chi^{\prime\prime}$. If $\gamma_{2}$ has length 2 or more, then $\chi^{\prime\prime}$ is nonempty, and we have $\alpha=ay(\chi^{\prime\prime}\eta\gamma_{1}a)\in L$ and $\alpha^{\prime}=(\gamma_{1}a)y\chi^{\prime\prime}\in L$. Let us take $y$ as body and consider the cross-product word $\alpha^{\prime\prime}=ay\chi^{\prime\prime}=a\gamma_{2}$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha\in L$, so as $L$ is reduced we know that $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged.
@@
Thanks to Section 6.1, we know that $\alpha=ax\eta ya\in L$ and $\alpha^{\prime}=yax\in L$, for some letters $x,y\in\Sigma$ and some word $\eta\in\Sigma^{*}$. So all that remains is to show:
{claim}
Let $L^{\prime}$ be a reduced language that contains words $ax\eta ya$ and $yax$ for some $a,x,y\in\Sigma$ and $\eta\in\Sigma^{*}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 12. Gadget used in the proof of Section 6.1
@@
{claim}
Let $L^{\prime}$ be a reduced language that contains a word of the form $aab$ for $a,b\in\Sigma$ with $a\neq b$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
@@
[…]
We can now conclude the proof of Section 6.1:
Figure 13. Gadget used in the proof of Section 6.1
Proof of Section 6.1.
We first rule out the case where $y=a$. Then $L^{\prime}$ contains $yax=aax$, and we conclude directly by Section 6.1 or Section 6.1 depending on whether $x=a$ or not.
We next rule out the case where $x=a$. Then, using Section 6.1, let us show hardness for $L^{\prime\mathrm{R}}$. This language contains $xay$, and as $x=a$ we can apply Section 6.1 or Section 6.1 again depending on whether $y=a$ or not, and conclude.
@@
To see why the condensed hypergraph of matches is as pictured on Figure 13, let us describe the directed paths in the completion of the gadget. The ones forming words of the form $ax\eta ya$, and $yax$, are guaranteed to be matches: the ones forming strict infixes of these are guaranteed not be matches because the language is reduced; and the ones containing one of these as an infix are guaranteed to be eliminated by node domination. In particular, paths that start within an $\eta$-edge, or end within an $\eta$-edge, or both, are not relevant. Indeed, observe that $\eta$-edges are always followed by (some prefix of) $yax\eta$. So a match starting within an $\eta$-edge (possibly using all facts of the edge) must include the subsequent $yax$ (because $\eta ya$ is a strict infix of $ax\eta ya$ and so is not in the language), and the match is then subsumed by the $yax$-match. Likewise, $\eta$-edges are always preceded by (some suffix of $\eta yax$), so matches ending within an $\eta$-edge (possibly using all facts of the edge) must include the preceding $yax$ (because $ax\eta$ is a strict infix of $ax\eta ya$) and then they are subsumed by the $yax$ match.
@@
For this reason, to build the condensed hypergraph of matches, it suffices to consider paths that start and end at the edges labeled $a$ and $x$ and $y$. One can check that such paths either are infixes of $ax\eta ya$ or contain $yax$ as an infix, so their membership to $L$ is determined by the fact that $L$ contains $yax$ and $ax\eta ya$ and is reduced. Building the corresponding condensed hypergraph of matches and further applying condensation rules gives Figure 13 (right), which is an odd chain, so we conclude by Section 4.8. ∎
@@
Applying Section 6.1 with $L^{\prime}\coloneq L$, we conclude that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard in the remaining cases of the non-overlapping case. This concludes the case distinction and concludes the proof of Section 6.1.
Connections to UCQs and self-joins.
Observe that finite languages $L$ correspond to the case of RPQs that are in fact unions of conjunctive queries (UCQs). The UCQs obtained from RPQs have a specific form: they are posed on an arity-two signature corresponding to the alphabet $\Sigma$, and their constituent conjunctive queries (CQs) are directed paths corresponding to words of the language $L$. In this sense, repeated letters in words correspond to self-joins in CQs. Thus, Section 6.1 implies that, for those UCQs that correspond to RPQs, the resilience problem is NP-hard unless every CQ is self-join-free. This contrasts with the setting of general UCQs, where resilience is sometimes tractable even in the presence of self-joins (DBLP:conf/pods/FreireGIM20). Specifically, in that setting, the following Boolean CQ on an arity-two signature is known to be in PTIME (DBLP:conf/pods/FreireGIM20) despite having two self-joins: $Q{\,:\!\!-\,}A(x),B(x),R(x,y),R(z,y),A(z),C(z)$.
Connections to other language classes.
We now explain how Section 6.1 relates to the languages discussed so far. We first perform a sanity check: such languages are never local.
{toappendix}
6.2. Proof of Section 6.2
{lemmarep}
{toappendix}
We point out that Section 6.2 does not in fact require the assumption that the language is reduced, and we do not use it in the proof:
Proof of Section 6.2.
As $L$ is finite and contains a word with a repeated letter, we can take a word with a repeated letter whose length is maximum among words of $L$ with a repeated letter. (Notice the similarity with the first criterion of the maximal-gap words of Definition 6.1.) Let $\alpha=\beta a\gamma a\delta\in L$ be such a word. Assume by contradiction that $L$ is local. Then, by Proposition 3.6, $L$ is letter-Cartesian. Applying the definition to $\alpha$ and $\alpha$, we must have that the following word is also in $L$:
$\beta a\gamma a\gamma a\delta\in L.$
Note that the finiteness hypothesis is essential, e.g., $L=ax^{*}b$ contains words with repeated letters but it is local (see Example 3.2). Of course it is also essential that $L$ is reduced, e.g., $L=a|aa$ contains a word with a repeated letter but its reduction $\mathtt{reduce}(L)=a$ is local.
Further notice that there is no converse to Section 6.2: there are finite reduced non-local languages that feature no words with repeated letters and for which resilience is hard. This is for instance the case of the four-legged language $axb|cxd$ of Proposition 4.8. Thus, the languages covered by Section 6.1 and Section 5.2 are incomparable: they each generalize Proposition 4.1 and Proposition 4.8, respectively, without implying the other result. (However, the proof of Section 6.1 relies on Section 5.2.) These two results also do not cover all intractable cases, as we will now see.

### v3

@@ 6. Hardness of Finite Languages Having Repeated Letters
First, we must have $\gamma_{1}\neq\epsilon$, otherwise $\alpha^{\prime}$ is an infix of $a\gamma$ which is a strict infix of $a\gamma a\delta\in L^{\prime}$, and this contradicts the assumption that $L^{\prime}$ is infix-free.
@@
Hence, applying 6.7 to $L^{\prime}\coloneq L$, we know that $L$ contains a maximal-gap word $\alpha=a\gamma a\delta$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}$ with $\gamma_{1}$ and $\gamma_{2}$ being respectively a non-empty suffix of $\gamma$ and a non-empty prefix of $\gamma$. We now claim that, if $\delta\neq\epsilon$, then $L$ is four-legged, allowing us to conclude:
Claim 6.8.
Let $L^{\prime}$ be an infix-free language containing a word $a\gamma a\delta$ such that $\delta\neq\epsilon$ and a word $\gamma_{1}a\gamma_{2}$ such that $\gamma_{1}\neq\epsilon$, $\gamma_{2}\neq\epsilon$, and $\gamma_{1}$ is a suffix of $\gamma$. Then $L^{\prime}$ is four-legged.
@@
Indeed, the word $\alpha^{\prime\prime}$ is an infix of $\gamma a\delta$ (because $\gamma_{1}$ is a suffix of $\gamma$), which is a strict infix of $a\gamma a\delta$ which is in $L^{\prime}$. Thus, as $L^{\prime}$ is infix-free, $\alpha^{\prime\prime}\notin L^{\prime}$, so we have a witness of the fact that $L^{\prime}$ is four-legged. ∎
Applying 6.8 with $L^{\prime}\coloneq L$, we can conclude unless $\delta=\epsilon$, which we assume in the rest of the proof.
@@
The overlapping case is when we have $|\gamma_{1}|+|\gamma_{2}|>|\gamma|$, so that we can write $\gamma=\eta^{\prime\prime}\eta\eta^{\prime}$ such that $\gamma_{1}=\eta\eta^{\prime}$ and $\gamma_{2}=\eta^{\prime\prime}\eta$, with the “overlap” $\eta$ being non-empty.
@@
Let us first deal with the overlapping case. We have $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, meaning $\alpha^{\prime}=\eta\eta^{\prime}a\eta^{\prime\prime}\eta\in L$, and we know that the “overlap” $\eta$ is non-empty. We then show:
Claim 6.9.
@@
Let us first assume that $\eta^{\prime}$ is non-empty and show that $L$ is four-legged in this case. Let us write $\eta^{\prime}=x\sigma$. Substituting this into $\alpha^{\prime}$ and $\alpha$, we have $\alpha^{\prime}=\eta x(\sigma a\eta^{\prime\prime}\eta)\in L$ and $\alpha=(a\eta^{\prime\prime}\eta)x(\sigma a)\in L$. Now, taking $x$ as body, let us consider the cross-product word $\alpha^{\prime\prime}=\eta x\sigma a=\eta\eta^{\prime}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta\eta^{\prime}a\in L$. Hence, as $L$ is infix-free, $\alpha^{\prime\prime}\notin L$, so that we can conclude that $L$ is four-legged unless $\eta^{\prime}$ is empty.
Let us substitute $\eta^{\prime}=\epsilon$ in $\alpha$ and $\alpha^{\prime}$. We know that $\alpha^{\prime}=\eta a\eta^{\prime\prime}\eta\in L$ and $\alpha=a\eta^{\prime\prime}\eta a\in L$. Let us now assume that $\eta^{\prime\prime}$ is non-empty and show that $L$ is four-legged in that case as well. Let us write $\eta^{\prime\prime}=x\sigma$. We have $\alpha=ax(\sigma\eta a)\in L$ and $\alpha^{\prime}=(\eta a)x(\sigma\eta)\in L$. Now, taking $x$ as body again, let us consider the cross-product word $\alpha^{\prime\prime}=ax\sigma\eta=a\eta^{\prime\prime}\eta$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\eta^{\prime\prime}\eta a\in L$. Hence, as $L$ is infix-free, $\alpha^{\prime\prime}\notin L$, so we can conclude again that $L$ is four-legged unless $\eta^{\prime\prime}$ is empty.
@@
From 6.9, we can conclude unless $\eta^{\prime}=\eta^{\prime\prime}=\epsilon$, so we assume this in the rest of the proof of hardness for the overlapping case. We now know that $\alpha^{\prime}=\eta a\eta\in L$ and $\alpha=a\eta a\in L$.
Let us now re-use the fact that $\alpha$ was chosen to be a maximal-gap word of $L$. Now, as $\eta$ is non-empty, observe that $\alpha^{\prime}=\eta a\eta$ contains repeated letters across the two copies of $\eta$. These letters are spaced at least by the length of $\eta$ (and possibly more, if $\eta$ contains repeated letters – even though this cannot happen by the maximal-gap condition). In any case, $\alpha^{\prime}$ achieves a gap between two repeated letters which is at least that of $\alpha$: so $\alpha$ and $\alpha^{\prime}$ are tied for the first criterion in the definition of maximal-gap. Hence, by the second criterion, since $\alpha$ is maximal-gap, we know that $\alpha^{\prime}$ cannot be strictly longer than $\alpha$. So from $|\alpha^{\prime}|\leq|\alpha|$, as $|\alpha^{\prime}|=2|\eta|+1$ and $|\alpha|=|\eta|+2$, we conclude $|\eta|\leq 1$. As $\eta$ is non-empty, we have $|\eta|=1$. This letter may be different from $a$ or not. So we can conclude the overlapping case by showing two hardness results:
Claim 6.10.
Let $L^{\prime}$ be any infix-free language containing abaaba and babbab. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is hard.
Figure 9. Completed gadget used in the proof of 6.10
@@
We use the gadget whose completion is pictured on Figure 9 (left). It is obviously a pre-gadget, and one can check that the hypergraph of matches can be reduced to the odd path of Figure 9 (right), so that it is indeed a gadget and we can conclude by Proposition 4.11. More precisely, observe that there are three kinds of directed walks in the gadget:
  • Paths of length 2 or less: all of these form a word which is a strict infix of abaaba or babbab, so they are not in the language by infix-freeness.
  • Paths of length 3: all of these form the word abaaba or babbab, so they are in the language.
@@
Hence, the hypergraph of matches of the language $L^{\prime}$ on the completion of the gadget corresponds precisely to the hyperedges given by the paths abaaba and babbab. One can then obtain the condensation of Figure 9 (right) by observing that $5\xrightarrow{b}1$ is subsumed by the next $a$-fact ($1\xrightarrow{a}2$) and that $8\xrightarrow{b}7$ is subsumed by the next $a$-fact ($7\xrightarrow{a}4$). The matches $\{1\xrightarrow{a}2,2\xrightarrow{b}3\}$ and $\{7\xrightarrow{a}4,4\xrightarrow{b}6\}$ then eliminate by edge domination the matches $\{t_{u}\xrightarrow{b}1,1\xrightarrow{a}2,2\xrightarrow{b}3\}$ and $\{1\xrightarrow{a}2,2\xrightarrow{b}3,3\xrightarrow{a}4\}$ and $\{t_{v}\xrightarrow{b}7,7\xrightarrow{a}4,4\xrightarrow{b}6\}$, and we can then see that $3\xrightarrow{a}4$ is eliminated by node domination (e.g., by $2\xrightarrow{b}3$) and that $t_{u}\xrightarrow{b}1$ is eliminated by node domination with $1\xrightarrow{a}2$ and $t_{v}\xrightarrow{b}7$ is eliminated by node domination with $7\xrightarrow{a}4$, leading to the pictured hypergraph. This allows us to conclude by Proposition 4.11. ∎
Claim 6.11.
Let $L$ be any infix-free language containing the word aaaaaa. Then $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard.
Figure 10. Completed gadget used in the proof of 6.11
@@
We use the gadget whose completion is pictured on Figure 10 (left) (which incidentally turns out to be identical to that of Figure 3(b) in Proposition 4.1). This is indeed a pre-gadget, and the hypergraph of matches can be condensed to Figure 10 (right), so it is indeed a gadget. Note that we use the fact that all directed walks in the gadget form words of the form $a^{*}$: by infix-freeness of the language, the matches precisely correspond to the words of length 3. ∎
@@
We conclude the proof of Theorem 6.1 by covering the non-overlapping case. Recall that, in this case, we know that the language $L$ contains a maximal-gap word $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$ and a word $\alpha^{\prime}=\gamma_{1}a\gamma_{2}\in L$, and recall that we already know that $\gamma_{1}$ and $\gamma_{2}$ are both non-empty.
@@
As $\gamma_{1}$ is non-empty, let us write $\gamma_{1}=\chi^{\prime}x$. If $\gamma_{1}$ has length 2 or more, then $\chi^{\prime}$ is non-empty, and we have $\alpha^{\prime}=\chi^{\prime}x(a\gamma_{2})\in L$ and $\alpha=(a\gamma_{2}\eta\chi^{\prime})xa\in L$. Let us take $x$ as body and consider the cross-product word $\alpha^{\prime\prime}=\chi^{\prime}xa=\gamma_{1}a$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha=a\gamma_{2}\eta\gamma_{1}a\in L$. So, as $L$ is infix-free, we have $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged. Hence, it suffices to focus on the case where $\gamma_{1}=x$ has length 1.
Now, as $\gamma_{2}$ is non-empty, let us write $\gamma_{2}=y\chi^{\prime\prime}$. If $\gamma_{2}$ has length 2 or more, then $\chi^{\prime\prime}$ is non-empty, and we have $\alpha=ay(\chi^{\prime\prime}\eta\gamma_{1}a)\in L$ and $\alpha^{\prime}=(\gamma_{1}a)y\chi^{\prime\prime}\in L$. Let us take $y$ as body and consider the cross-product word $\alpha^{\prime\prime}=ay\chi^{\prime\prime}=a\gamma_{2}$. The word $\alpha^{\prime\prime}$ is a strict infix of $\alpha\in L$, so as $L$ is infix-free we know that $\alpha^{\prime\prime}\notin L$, witnessing that $L$ is four-legged.
@@
Thanks to 6.12, we know that $\alpha=ax\eta ya\in L$ and $\alpha^{\prime}=yax\in L$, for some letters $x,y\in\Sigma$ and some word $\eta\in\Sigma^{*}$. So all that remains is to show:
Claim 6.13.
Let $L^{\prime}$ be an infix-free language that contains words $ax\eta ya$ and yaxyax for some $a,x,y\in\Sigma$ and $\eta\in\Sigma^{*}$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
Figure 11. Completed gadget used in the proof of 6.14
@@
Claim 6.14.
Let $L^{\prime}$ be an infix-free language that contains a word of the form aabaab for $a,b\in\Sigma$ with $a\neq b$. Then $\mathtt{RES}_{\mathrm{set}}(L^{\prime})$ is NP-hard.
@@
We use the gadget whose completion is pictured on Figure 11 (left). Let us now discuss the condensed hypergraph of matches. First observe that there are three matches forming the word aabaab, and walks of length at most 2 must form strict infixes of aabaab so they do not correspond to words of $L^{\prime}$. Note that, as $ab$ is not in $L^{\prime}$ (as a strict infix of $aab\in L^{\prime}$), then we know that every match using $3\xrightarrow{a}1$ must also use the fact $t_{v}\xrightarrow{a}3$; hence $3\xrightarrow{a}1$ can be eliminated by node-domination. Likewise, $b$ is not in $L^{\prime}$, so every match using $3\xrightarrow{b}4$ must also use $t_{v}\xrightarrow{a}3$, and $3\xrightarrow{b}4$ is eliminated by node domination. This ensures that the only walk of length greater than 3 which does not contain aabaab as an infix, namely, the aaaaaa path from $s_{v}$ to $1$, is subsumed in the condensation in the case where $aaa\in L^{\prime}$: namely, after eliminating $3\xrightarrow{a}1$ by node elimination, the remaining facts $\{s_{v}\xrightarrow{a}t_{v},t_{v}\xrightarrow{a}3\}$ correspond to the match $\{s_{v}\xrightarrow{a}t_{v},t_{v}\xrightarrow{a}3,3\xrightarrow{b}4\}$ after we have used node domination to eliminate $3\xrightarrow{b}4$. Thus, we can build a condensed hypergraph of matches by considering only the walks that form the word aabaab, and then applying condensation we obtain Figure 11 (right). This establishes that we have a gadget, so we can conclude by Proposition 4.11. ∎
We can now conclude the proof of 6.13:
Figure 12. Completed gadget used in the proof of 6.13
Proof of 6.13.
We first rule out the case where $y=a$. Then $L^{\prime}$ contains $yax=aax$, and we conclude directly by 6.14 or 6.11 depending on whether $x=a$ or not.
We next rule out the case where $x=a$. Then, using Proposition 6.3, let us show hardness for $L^{\prime\mathrm{R}}$. This language contains xayxay, and as $x=a$ we can apply 6.14 or 6.11 again depending on whether $y=a$ or not, and conclude.
@@
To see why the condensed hypergraph of matches is as pictured on Figure 12, let us describe the directed walks in the completion of the gadget. The ones forming words of the form $ax\eta ya$, and yaxyax, are guaranteed to be matches: the ones forming strict infixes of these are guaranteed not be matches because the language is infix-free; and the ones containing one of these as an infix are guaranteed to be eliminated by node domination. In particular, walks that start within an $\eta$-edge, or end within an $\eta$-edge, or both, are not relevant. Indeed, observe that $\eta$-edges are always followed by (some prefix of) $yax\eta$. So a match starting within an $\eta$-edge (possibly using all facts of the edge) must include the subsequent yaxyax (because $\eta ya$ is a strict infix of $ax\eta ya$ and so is not in the language), and the match is then subsumed by the yaxyax-match. Likewise, $\eta$-edges are always preceded by (some suffix of $\eta yax$), so matches ending within an $\eta$-edge (possibly using all facts of the edge) must include the preceding yaxyax (because $ax\eta$ is a strict infix of $ax\eta ya$) and then they are subsumed by the yaxyax match.
@@
For this reason, to build the condensed hypergraph of matches, it suffices to consider walks that start and end at the edges labeled $a$ and $x$ and $y$. One can check that such walks either are infixes of $ax\eta ya$ or contain yaxyax as an infix, so their membership to $L$ is determined by the fact that $L$ contains yaxyax and $ax\eta ya$ and is infix-free. Building the corresponding condensed hypergraph of matches and further applying condensation rules gives Figure 12 (right), which is an odd path, so we conclude by Proposition 4.11. ∎
@@
Applying 6.13 with $L^{\prime}\coloneq L$, we conclude that $\mathtt{RES}_{\mathrm{set}}(L)$ is NP-hard in the remaining cases of the non-overlapping case. This concludes the case distinction and concludes the proof of Theorem 6.1.